# Quantum Hall and Light Responses in a 2D Topological Semimetal

### 06 Dec 2023 v2

@@ Abstract
We investigate the topological characteristics of a recently discovered class of semimetals in two dimensions on the honeycomb lattice. These semimetals reside at the transition between two distinct topological insulators, each existing in a nontrivial topological phase. As a result, these semimetals exhibit specific topological properties, including the presence of edge modes. In a preceding work, we demonstrated the topological robustness of this semimetal phase against disorder and interactions. In this work, we delve deeper into the semimetal’s electronic properties, providing a precise calculation of its Hall conductivity and response to circularly polarized light, elaborating further on its bulk-edge correspondance leading to a half topological semimetal.
@@ I Introduction
Topological materials have emerged in the last decades as a remarkable class of materials that challenge conventional classifications of matter [1]. Among these materials, two-dimensional topological systems display unique electronic properties such as a quantized Hall conductivity [2, 3]. Moreover, unlike conventional insulators, which have an energy gap between their valence and conduction bands throughout the entire material, these quantum Hall-like systems possess states that display metallic behavior [4, 5] which are exclusively located on the edges allowing for the current to flow around the sample, protected from impurities and disorder.
@@
At the heart of topological aspects in these condensed-matter systems is the nontrivial nature of the wavefunction in reciprocal space associated to the band structure of a crystal. Indeed, one of the key theoretical concepts underlying the study of topological matter is the notion of topological invariants which have direct consequences on the dynamics of the material. The main topological invariant of interest is the Chern number associated to the Brillouin Zone of the material. Remarkably, this integer Chern number was found to be directly related to the Hall conductivity of the material, showing a unique quantized behavior in units of the fundamental constant $e^{2}/h$ [3, 6], accessible in quantum transport experiments. This topological number was also found to be equivalently accessible from the response of the material to circularly polarized light [7, 8, 9]. This topological response to circularly polarized light can also be equivalently resolved locally within the Brillouin zone [10].
Figure 1: Color online: (Above) One-dimensional energy bands of the Kane-Mele model in the quantum spin Hall phase for the two spin polarizations of an electron, related through time-reversal symmetry. The spin polarization is shown in the third direction. (Bottom) Resulting energy bands in the momentum-Energy axis. Energies are measured in units of $t$ for the different figures.
@@
Quantum spin Hall insulators are two-dimensional topological insulators which have gained prominence in the last two decades [11, 12]. Quantum spin Hall insulators are characterized by the presence of the spin Hall effect, where an electric current induces a transverse spin current due to intrinsic spin-orbit coupling. These spin Hall systems are most often theoretically described by two spin-polarized subsystems, called spin channels, which are related through time-reversal symmetry. A specific example is shown in Fig. 1 for the case of the Kane-Mele model [13, 14] in which the spin of the electron is a well defined quantum number labelling distinct energy bands. In this model, each spin channel hosts oppositely directed edge modes, referring then to the quantum spin Hall effect. This description in terms of counter-propagating spin channels is also valid for the Bernevig-Hughes-Zhang model [15]. The system then hosts a zero total Chern number and is equivalently described through a $\mathbb{Z}_{2}$ topological spin Chern number related to the structure of the quantum spin Hall current at the edges [13, 16]. Circularly polarized light can measure this topological $\mathbb{Z}_{2}$ invariant related to the quantum spin Hall effect [10] and the parity symmetry [17] with a perfect quantization at the Dirac points on the honeycomb lattice associated to the zeros of the Pfaffian [14].
The classification of topological properties in terms of two distinct Pauli matrices, e.g. related to the two sublattices on the honeycomb lattice and to the two spin polarizations of an electron in the Kane-Mele model, was at the heart of our recent publication [18] in which we describe a possible generalization of nontrivial topological properties and robustness towards a specific class of two-dimensional topological semimetals. The first hints pointing at the topological nature of this semimetal came from an earlier study made in our group on the general framework of topological spheres’ models [19], and in which this topological semimetal was found to sit at a transition between two topologically distinct but still non trivial phases of the system. This semimetal was then understood to host two independent copies of the Haldane model [20] as in the Kane-Mele model, however in this case the two copies hosted the same topology but each was partially filled differently. As such, this semimetal was found to host robust edge modes characteristics of topological materials, giving rise to its non trivial topological nature as a semimetal. The robust edge mode is polarized along $x$ direction which has then a one-half probability to reside in each plane for the bilayer model [19] or to be polarized along $z$ direction for the one-layer graphene model [18].
The question of the topological classification of the semimetals is still an open question to this day [21], related to a possible generalization of the Thouless-Kohmoto-Nightingale-denNijs (TKNN) formula [3] for systems with a Fermi surface [22]. Indeed, for topological insulators, this formula expresses a dynamical property of the system (its Hall conductance) as a function of a topological property of the wavefunction of the ground state (its Chern number in reciprocal space). Several schemes can then be adopted in order to find a topological classification for semimetals, taking an inspiration from the existing classification for insulators. A first strategy would be to investigate in the gapless band structure of the semimetal possible candidates for topological invariants [23] but these invariants may not correspond to directly accessible experimental observables in terms of electronic transport. A second approach would be to discuss the topological properties of the corresponding insulators if the gap of the concerned semimetal opened and made infinitesimally small [24, 25] opening this way the field of approximate symmetries. A third approach, as done in the present paper, would consist to investigate in the band structure of the semimetal possible signatures of the topological nature of the bands. Yet, taking the point of view of the first approach, we will show how the protocols of [9, 10, 26] to analyze the response to circularly polarized light locally within the Brillouin zone can also provide a meaningful information of the topological response of the quantum Hall conductivity related to a half topological semimetal [19, 27]. This approach is restricted here to a specific class of models which are built out of two or more independent copies of the Haldane model, but this class of models is still of major interest nowadays for this discussion to bring many interesting insights to the field of topology in condensed matter physics.
In this context, the aim of the present paper is thus to provide a clear evaluation of the quantum Hall conductivity and of the responses to circularly polarized light for this protected topological semimetal discussing in greater details its bulk-edge correspondance. The article is organized as follows. In Sec. II, we remind main properties of the topological semimetal. In Sec. III, we present the derivation of the quantum Hall conductivity for this system. In Sec. IV, we derive a topological response through circularly polarized light locally from the Dirac points within the Brillouin zone. In Sec. V, we discuss the bulk-edge correspondence and possible applications through Angle-Resolved Photoemission Electron Spectroscopy (ARPES), and formulate general remarks on this theoretical scheme. Then, in Sec. VI, we summarize our findings.
@@ II Protected topological semimetal
The model of the semimetal we are interested in can be realized in a single layer of graphene [18]. This model relies on the presence of an Haldane model associated to the two spin polarizations of an electron. We include the presence of a magnetic field, such that its direction is fixed parallel to the layer to induce Zeeman effects. The energy bands of the system will be then classified according to the two spin polarizations of an electron along $x$ direction. In the recent years, there have been many developments of proximity-induced topological phases of matter [28, 29, 30, 31] and in particular on graphene/Transition-metal dichalcogenide (TMD) heterostructures [32, 33] with easily tunable parameters. Therefore, we can anticipate that there will be various equivalent ways to realize this topological semimetal in two dimensions through two sets of Pauli matrices related to the physics of heterostructures. The Hamiltonian of the two-dimensional (2D) topological semimetal can be written in tight-binding formalism in reciprocal space as [18]
@@
Figure 2: Color Online: Two-dimensional band structure of the system for a chosen path in the Brillouin zone with the spin polarizations associated to the energy bands shown in distinct colors. The parameters chosen for all panels are $M=3\sqrt{3}t/4$, $r=t/\sqrt{3}$, $t_{2}=t/3$ with lattice spacing $a=1$. Results can be generalized as long as Eq. (11) is satisfied.
@@
[…]
$3\sqrt{3}t_{2}-M\leq\tilde{r}<3\sqrt{3}t_{2}+M.$ (11)
The parameter $\tilde{r}$ can be viewed as a small renormalization effect of the transverse magnetic field as a result of interactions which can be elegantly taken care of through a variational stochastic approach which can be successfully applied to interacting topological models [9, 36, 37] in quantitative agreement with numerical methods [38, 39]. Hereafter, we investigate in further details the expression of the Hall conductivity of the material to see how this can be related to topological invariants. This computation is in general not possible for crossing bands but we argue below that for this specific model, the evaluation is well defined and is free of divergences e.g. as a result of orthogonality of spin eigenstates on both sides of the crossing regions.
For the Haldane model, or generally speaking a topological two-bands model, the responses to circularly polarized light reveal the same information as the quantum Hall conductivity [7, 8, 9] and this information can be in fact re-built from specific points in the Brillouin zone [10, 26], in particular the Dirac points.
We will then address this correspondence between quantum Hall conductivity and responses to circularly polarized light in Sec. IV and show that this yet reveals important signatures of topological properties of the semimetal from the Dirac points, e.g. through the topological invariant of the lowest-energy band, with a correspondence onto a half topological semimetal in the basis of $\uparrow_{z},\downarrow_{z}$ spin states.
III Kubo formula for Hall conductivity
In the linear response regime, the Kubo formula for the Hall conductivity of a 2D electron gas is given by [3]
@@ III Kubo formula for Hall conductivity
One may question the applicability of this formula in the semimetallic regime for which bands $\downarrow_{x},+$ and $\uparrow_{x},-$ cross zero energy such that $E_{n}(\bm{k})-E_{m}(\bm{k})=0$, which seems to give us a diverging fraction in the expression of Eq. (III). Fortunately, this divergence is avoided by the fact that the current operator $v^{\alpha}$ does not couple these two bands as we have
@@
such that $\langle\uparrow_{x},-|v^{\alpha}|\downarrow_{x},+\rangle=0$ identically over the entire Brillouin zone. Eq. (III) is thus well defined and we can rewrite this expression in terms of the integral of the Berry curvature over the Brillouin zone over all occupied states
@@
In the semimetallic context, the contribution from the Fermi-Dirac distribution with the term $f(E_{n}(\bm{k}))$ appearing in Eq. (14) plays a major role as it constrains the integration to be performed only partially over the Brillouin zone for the two middle bands $n=\downarrow_{x},+$ and $n=\uparrow_{x},-$ which correspond to two partially filled bands, while the integration is performed over the entire Brillouin zone for $n=\downarrow_{x},-$ whose band is entirely filled. This expression gives
@@
As discussed by Haldane in 2004 [22], the partial integrals can still be seen as the Berry flux of particles moving around the closed Fermi surface for each band and can be interpreted as a topological property of the Fermi surface when looking to the semimetal as a Fermi liquid [18], but these integrals are not topological invariants.
Figure 3: Numerical computation of the Hall conductivity of the topological semimetal for a given sample using the Kubo formula, as shown in Eq. (III). The total Hall conductivity gradually decreases from its quantized value $|\sigma^{xy}|=2e^{2}/h$ in the topological insulating phase to $|\sigma^{xy}|=0$ in the trivial insulating phase according to Eq. (32).
Furthermore, the $\bm{d}_{\bm{k}}$ vector appearing in the Hamiltonian of Eq. (3) gives us the following well defined mapping from the Brillouin Zone to 2-sphere
$\Phi:BZ$ $\rightarrow S^{2}$ (20)
@@
We can then integrate formally by change of variables $\bm{k}\rightarrow(\theta,\varphi)$. The Fermi surface formed by the nodal ring can be expressed to a good approximation in these variables by the contour $\theta=\theta_{c},\varphi\in\left[0,2\pi\right]$ for a given $\theta_{c}$ fixed between 00 and $\pi$ depending on the value of $r$, which defines the position of the nodal ring in the Brillouin zone.
@@
We see that this expression interpolates between $\sigma^{xy}=-2\dfrac{e^{2}}{h}$ for the topological insulating phase when $\tilde{r}<3\sqrt{3}t_{2}-M$ for which $\theta_{c}=\pi$ all the way to $\sigma^{xy}=0$ for the trivial insulating phase when $\tilde{r}>3\sqrt{3}t_{2}+M$ for which $\theta_{c}=0$. The nodal ring equivalently navigates from the vicinity of $K^{\prime}$ towards $K$ when increasing $\tilde{r}$ within the semimetal region such that $3\sqrt{3}t_{2}-M<\tilde{r}<3\sqrt{3}t_{2}+M$. These results do agree with the numerical analysis shown in Fig. 3 which represents the numerical evaluation of Eq. (III). The total Hall conductivity is thus no longer proportional to specific topological invariants in the topological semimetallic phase. However, as shown in Eq. (16), its value can be decomposed into several contributions each having its own topological interpretation, in terms of topological invariant for the $C_{\downarrow_{x},-}$ term, or in terms of Berry flux accumulated by particles moving around the Fermi surface for the two other terms [22].
Below, we show that the quantized topological response of the lowest band associated to $C_{\downarrow_{x},-}$ can be measured through circularly polarized light related to a half topological semimetal. Then, we will address the open question of the bulk-edge correspondance in the topological semimetal.
IV Circularly Polarized Light
Here, we describe the effect of a circularly polarized light field in the plane [7, 9, 26, 10].
We begin with one plane described through a Haldane model within the topological phase to fix the definitions. Within the Dirac approximation, we take the same definitions of the Brillouin zone as in the articles [9, 26], such that the current operator takes the form
@@ IV Circularly Polarized Light
with the velocity $v_{F}=\frac{3}{2\hbar}ta$ and ${\bf p}=(p_{x},p_{y})$ measures a small deviation from the Dirac points. Here, $\zeta=\pm 1$ at the $K$ and $K^{\prime}$ respectively such that within the choice of the Brillouin zone the two Dirac points are related through $p_{y}\rightarrow-p_{y}$ or more generally $k_{y}\rightarrow-k_{y}$. The vector potential due to the classical light source is included through a shift of the momentum. There is a subtlety to mention here within this definition: $\zeta$ refers both to the direction of light polarization, right-handed (+) or left-handed (-), and to the Dirac points of interest such that
@@
For the Haldane model, this leads to the following transition rates related to the inter-band light-induced transition probabilities [9, 10, 26]
@@
An important point is to incorporate the energy conservation within the four-energy bands model. It is then useful to write down the light-matter interaction $\delta H_{\pm}=A_{0}e^{\pm i\omega t}\sigma^{+}\otimes\mathbb{I}$ within the four energy bands. Since the light-matter interaction couples bands with the same spin quantum number this implies that at the $K$-point this will only mediate transitions between bands $n=\downarrow_{x},-$ and band $m=\downarrow_{x},+$ and separately between bands $n^{\prime}=\uparrow_{x},-$ and band $m^{\prime}=\uparrow_{x},+$. Interestingly, it is possible to find a resonance with both pairs of bands fixing the light frequency such that $\hbar\omega=2|d_{z}(K)|=6\sqrt{3}t_{2}+2M$ for the right-handed wave (+). It is interesting to observe that in the present case, it is possible to fix the resonant condition just with the $K$ Dirac point due to the precise shape of the energy bands in Fig. 1.
@@
A peculiarity of this topological semimetal is that $\langle\sigma_{z}(\pi)\rangle=0$ as a result of the formation of the nodal ring semimetal around $K^{\prime}$ [19] which is another way to justify why $\Gamma_{-}=0$ in this case. The response for one-spin polarization is now halved compared to one topological plane as a result of $\Gamma_{-}=0$. The topological semimetal can be characterized in terms of a pair of one-half local topological markers and $\pi$-Berry phases [27] related to the two-spheres’ model associated to spin polarizations $\uparrow_{z},\downarrow_{z}$ [19]. In the present situation, the $\pi$ Berry phase is associated equally to the phase acquired by an electron with spin polarization $\uparrow_{x}$ in band $2$ and by an electron with spin polarization $\downarrow_{x}$ in band $1$, when circling around the $K$ Dirac. A pair of $\pi$-Berry phases is also analogous to a pair of $\pi$ winding numbers [27]. A $\frac{1}{2}$ topological quantized light response is also predicted to occur for the $\nu=\frac{1}{2}$-fractional quantum Hall effect [42].
@@
Therefore, we obtain a relation between light and quantum Hall response for the topological semimetal
@@
We have reminded above that the light response at the $K$ and $K^{\prime}$ points for the topological Haldane model reveals the same information as the light response and therefore the quantum Hall response coming from a summation on the whole Brillouin zone [26]. This identification necessitates that the light response functions are continuous and differentiable everywhere. For the topological semimetal, for the intermediate energy bands, the function $F_{p_{x}p_{y}}$ flips its sign at $\theta_{c}$ related to the analysis of Sec. III. In this way, the light response from the Dirac points is identical to the quantum Hall response of the semimetal only if $F_{p_{x}p_{y}}=0$ at $\theta_{c}$ for the second band. In the region $\overline{NR}$, the function $F_{p_{x}p_{y}}$ is identical for the two lowest energy bands. Therefore, we must set $\theta_{c}=\frac{\pi}{2}$ to be sure that the light response from the summation on the whole Brillouin zone would be equivalent to the light responses at the Dirac points. Fixing $\theta_{c}=\frac{\pi}{2}$, the conductivity response from the second-energy band becomes also zero. This is then another way to justify that the light response at the Dirac points reveals the quantum Hall conductivity of the lowest energy band and therefore the topological invariant, which remains identical for all the range of parameters in Eq. (11).
For $\theta_{c}=\frac{\pi}{2}$, we also observe that the quantum Hall response coming from $\uparrow_{x}$ particles in the second-energy band takes the form $\frac{e^{2}}{h}\widehat{C}_{\uparrow_{x},-}=-\frac{1}{2}\frac{e^{2}}{h}$ which justifies further the introduction of a half topological number for this spin polarization [27]. Similarly, if we now sum the topological number associated to the $\downarrow_{x}$ particles in the two energy bands, this also measures $\frac{e^{2}}{h}(\widetilde{C}_{\downarrow_{x},+}+C_{\downarrow_{x},-})=-\frac{1}{2}\frac{e^{2}}{h}$ and gives a similar interpretation to the other half-topological number. This approach then justifies further the relation between the topological response of the quantum Hall conductivity and $\sum_{j=\uparrow,\downarrow}C_{\alpha}$ with $|C_{\alpha}|=\frac{1}{2}$ mentioned in previous publications [19, 18, 27].
When $\tilde{r}<3\sqrt{3}t_{2}-M$, the same light polarization would equally interact with both pairs $(n,m)$ and $(n^{\prime},m^{\prime})$ at one Dirac point measuring twice the topological invariant of the lowest band, in agreement with the quantum Hall conductivity. For $\tilde{r}>3\sqrt{3}t_{2}-M$, one pair of bands will interact with one distinct light polarization at each Dirac point. In this case, if we reproduce the steps of the calculation above we now find that $\Gamma_{+}-\Gamma_{-}=0$ at each Dirac point. In Eq. (36), the two pairs $(n,m)$ and $(n^{\prime},m^{\prime})$ are now characterized with different signs of $A_{y}(t)$ and inverted lower and upper energy eigenstates, for the same Dirac point.

### 27 Feb 2024 v3

@@ Abstract
We have recently identified a protected topological semimetal in graphene which presents a zero-energy edge mode robust to disorder and interactions. Here, we address the characteristics of this semimetal and show that the $\mathbb{Z}$ topological invariant of the Hall conductivity associated to the lowest energy band can be equivalently measured from the resonant response to circularly polarized light resolved at the Dirac points. The (non-quantized) conductivity responses of the intermediate energy bands, including the Fermi surface, also give rise to a $\mathbb{Z}_{2}$ invariant. We emphasize on the bulk-edge correspondence as a half-topological protected semimetal, i.e. one spin-population polarized in the plane is in the insulating phase related to the robust edge mode while the other is in the metallic regime. The quantized transport at the edges is also equivalent to a $\frac{1}{2}-\frac{1}{2}$ conductance for spin polarizations along $z$ direction. We also build a parallel between the topological Hall response through the quantized light response at the Dirac points and a pair of half numbers (half Skyrmions).
@@ I Introduction
At the heart of topological aspects in condensed-matter systems is the nontrivial nature of the wavefunction in reciprocal space associated to the band structure of a crystal. Indeed, one of the key theoretical tools underlying the study of topological matter is the notion of topological invariants which have direct consequences on the dynamics of the material. The main topological invariant of interest is the Chern number associated to the Brillouin Zone of the material. Remarkably, this integer Chern number was found to be directly related to the Hall conductivity of the material, showing a unique quantized behavior in units of the fundamental constant $e^{2}/h$ [1, 2], accessible in quantum transport experiments. This topological number was also found to be equivalently accessible from the response of the material to circularly polarized light [3, 4, 5]. This topological response to circularly polarized light can also be equivalently resolved locally within the Brillouin zone [6]. The topological characterization can also be revealed from a local microscope i.e. a long transmission line transporting classical light [7].
@@
Quantum spin Hall insulators are two-dimensional topological insulators which have gained prominence in the last two decades [8, 9]. Quantum spin Hall insulators are characterized by the presence of the spin Hall effect, where an electric current induces a transverse spin current due to intrinsic spin-orbit coupling. These spin Hall systems are most often theoretically described by two spin-polarized subsystems, called spin channels, which are related through time-reversal symmetry. A specific example is the case of the Kane-Mele model [10, 11, 12, 13, 14] in which the spin of the electron is a well defined quantum number labelling distinct energy bands. In this model, each spin channel hosts oppositely directed edge modes, referring then to the quantum spin Hall effect. This description in terms of counter-propagating spin channels is also valid for the Bernevig-Hughes-Zhang model [15]. The system then hosts a zero total Chern number and is equivalently described through a $\mathbb{Z}_{2}$ topological spin Chern number related to the structure of the quantum spin Hall current at the edges [10, 16]. Circularly polarized light can measure this topological $\mathbb{Z}_{2}$ invariant related to the quantum spin Hall effect [6] and the parity symmetry [17] with a perfect quantization at the Dirac points on the honeycomb lattice associated to the zeros of the Pfaffian [11].
The classification of topological properties in terms of two distinct Pauli matrices, e.g. related to the two sublattices on the honeycomb lattice and to the two spin polarizations of an electron in the Kane-Mele model, was at the heart of our recent publication [18], in which we describe a possible generalization of nontrivial topological properties and robustness towards a specific class of two-dimensional (2D) topological semimetals realizable in graphene. The topological protection of this semimetal was shown through a robust zero-energy edge mode polarized along $x$ direction in the energy spectrum. This was shown numerically in the presence of disorder and through a stochastic variational approach in the presence of interactions [5, 19, 20].
In this article, we study the bulk topological characteristics of this 2D semimetal in graphene and we present a precise correspondence with the quantum Hall conductivity and the response to circularly polarized light. In Sec. II, we recall the Hamiltonian and band structure of this system [18]. Applying the Kubo formula for the topological semimetal, in Sec. III, we check that the topological invariant of the lowest-energy band spin-polarized along $x$ direction remains quantized for all the range of parameters in the sense of the quantum Hall response, revealing the existence of the edge mode. An interesting result found in Sec. IV is that the response to circularly polarized light resolved in frequency, corresponding to a resonance with the Dirac points, can equally reveal this quantized $\mathbb{Z}$ topological marker associated to the lowest-energy band. In Sec. III we also derive the quantum Hall conductivity of the second-energy energy band and of the nodal ring region. We verify that these contributions are well defined, but these integrals of Berry flux of particles are not quantized reflecting the Fermi surface properties in agreement with the article by Haldane in 2004 [21]. Yet, we find that we can introduce a quantized $\mathbb{Z}_{2}$ topological marker to characterize the topological properties associated to the second-energy band and the Fermi surface. The light response at the Dirac points for the semimetal is equivalent to a pair of $\frac{1}{2}$ topological numbers (half Skyrmions), linking then with the physics of the topological semimetal in the bilayer model related to the model of two entangled spheres [22, 23]. This semimetal was understood to host two channels of the Haldane model [24] as in the Kane-Mele model, revealing a $\mathbb{Z}_{2}$ symmetry and halved transport at the edges. In the present study in one layer graphene, the bulk-edge correspondence in Sec. V will lead to the formation of a half topological semimetal when varying the chemical potential i.e. with one spin polarization showing a quantized conductance and one spin polarization showing a metallic behavior. The quantized transport at the edges along $x$ direction is also equivalent to a $\frac{1}{2}-\frac{1}{2}$ probability for each spin polarization along $z$ direction showing a correspondence with the topological bilayer system [22, 23].
@@ II Protected topological semimetal
The model of the semimetal in one layer graphene [18] is related to the presence of an Haldane model associated to the two spin polarizations of an electron in the plane. We include a magnetic field, such that its direction is fixed parallel to the layer to induce Zeeman effects. The energy bands of the system will be then classified according to the two spin polarizations of an electron along $x$ direction. In recent years, there have been many developments of proximity-induced topological phases of matter [25, 26, 27, 28] and in particular on graphene/Transition-metal dichalcogenide (TMD) heterostructures [29, 30] with easily tunable parameters. Therefore, we can anticipate that there will be various equivalent ways to realize this topological semimetal in two dimensions through two sets of Pauli matrices related to the physics of heterostructures. The Hamiltonian of the two-dimensional (2D) topological semimetal can be written in tight-binding formalism in reciprocal space as [18]
@@
Figure 1: Color Online: Two-dimensional band structure of the system for a chosen path in the Brillouin zone with the spin polarizations associated to the energy bands shown in distinct colors. The parameters chosen for all panels are $M=3\sqrt{3}t/4$, $r=t/\sqrt{3}$, $t_{2}=t/3$ with lattice spacing $a=1$.
@@
III Quantum Hall Response and Invariants
In the linear response regime, the Kubo formula for the Hall conductivity of a 2D electron gas reads [1]
@@ III Quantum Hall Response and Invariants
One may question the applicability of this formula in the semimetallic regime for which bands $\downarrow_{x},+$ and $\uparrow_{x},-$ cross zero energy such that $E_{n}(\bm{k})-E_{m}(\bm{k})=0$. Here, a divergence does not occur because the current operator
@@
shows the important property $\langle\uparrow_{x},-|v^{\alpha}|\downarrow_{x},+\rangle=0$. Eq. (III) is thus well defined and we can rewrite this expression in terms of the integral of the Berry curvature on the Brillouin zone and all occupied states
@@
In the semimetallic regime, the contribution from the Fermi-Dirac distribution with the term $f(E_{n}(\bm{k}))$ appearing in Eq. (13) plays a major role as it constrains the integration to be performed only partially over the Brillouin zone for the two middle bands $n=\downarrow_{x},+$ and $n=\uparrow_{x},-$ which correspond to two partially filled bands, while the integration is performed over the entire Brillouin zone for $n=\downarrow_{x},-$ whose band is entirely filled. This expression gives
@@
From a mathematical point of view, it is useful to formulate a correspondence from the Brillouin zone to 2-sphere. The same map will also clarify the response to circularly polarized light in the next Section. The $\bm{d}_{\bm{k}}$ vector appearing in the Hamiltonian of Eq. (3) gives us the following well defined mapping from the Brillouin Zone to 2-sphere
$\Phi:BZ$ $\rightarrow S^{2}$
@@
We can then integrate by change of variables $\bm{k}\rightarrow(\theta,\varphi)$. The Fermi surface formed by the nodal ring can be expressed to a good approximation in these variables by the contour $\theta=\theta_{c},\varphi\in\left[0,2\pi\right]$ for a given $\theta_{c}$ fixed between 00 and $\pi$ depending on the value of $r$, which defines the position of the nodal ring in the Brillouin zone.
@@
This expression interpolates between $\sigma^{xy}=-2\dfrac{e^{2}}{h}$ for the topological insulating phase when $r<3\sqrt{3}t_{2}-M$ for which $\theta_{c}=\pi$ all the way to $\sigma^{xy}=0$ for the trivial insulating phase when $r>3\sqrt{3}t_{2}+M$ for which $\theta_{c}=0$. The nodal ring equivalently navigates from the vicinity of $K^{\prime}$ towards $K$ when increasing $r$ within the semimetal region such that $3\sqrt{3}t_{2}-M<r<3\sqrt{3}t_{2}+M$. However, the quantized invariant for the $C_{\downarrow_{x},-}$ term associated to the lowest-energy band characterizes the presence and robustness of the zero-energy edge mode and can therefore represent a good marker for this topological system. Now, regarding the other two terms traducing the Berry flux accumulated by particles moving around the Fermi surface [21], we can in fact also introduce a quantized $\mathbb{Z}_{2}$ invariant, valid for all range of parameters within the semimetal metal phase, such that
$\hat{C}_{\uparrow_{x},-}-\tilde{C}_{\downarrow_{x},+}=C_{\downarrow_{x},-}.$ (31)
In this sense, we can introduce two invariants to characterize this energy band structure i.e. the $\mathbb{Z}$ quantized invariant related to the lowest energy band and a $\mathbb{Z}_{2}$ index formed through the second-energy band and the nodal ring region or Fermi surface.
Below, we show that the quantized topological response of the lowest band associated to $C_{\downarrow_{x},-}$ can be measured through circularly polarized light which then reveals the presence of the robust edge mode. We will also show that this response locally measured from the resonance at the Dirac points is mathematically equivalent to fix $\theta_{c}=\frac{\pi}{2}$ on a sphere i.e. revealing a half-Skyrmion, linking with the physics of the bilayer model [22, 23].
IV Topological Invariants from Circularly Polarized Light
Here, we measure the topological response from a circularly polarized light field in the plane [3, 5, 31, 6] localized resolved in frequency to reach the Dirac point(s). We begin with one plane described through a Haldane model within the topological phase to fix the definitions. Within the Dirac approximation, we take the same definitions of the Brillouin zone as in [5, 31], such that the (photo-induced) current operator takes the form,
@@ IV Topological Invariants from Circularly Polarized Light
with the velocity $v_{F}=\frac{3}{2\hbar}ta$ and ${\bf p}=(p_{x},p_{y})$ measures a small deviation from the Dirac points. Here, $\zeta=\pm 1$ at the $K$ and $K^{\prime}$ respectively such that within the choice of the Brillouin zone the two Dirac points are related through $p_{y}\rightarrow-p_{y}$ or more generally $k_{y}\rightarrow-k_{y}$. The vector potential due to the classical light source is included through a shift of the momentum. There is a subtlety to mention here: $\zeta$ refers both to the direction of light polarization, right-handed (+) or left-handed (-), and to the Dirac points of interest such that
@@
For the Haldane model, this leads to the following response associated to the photo-induced currents [5, 31]
@@
An important point is to incorporate the energy conservation within the four-energy bands model. It is then useful to write down the light-matter interaction $\delta H_{\pm}=A_{0}e^{\pm i\omega t}\sigma^{+}\otimes\mathbb{I}$ within the four energy bands. Since the light-matter interaction couples bands with the same spin quantum number this implies that at the $K$-point this will only mediate transitions between bands $n=\downarrow_{x},-$ and band $m=\downarrow_{x},+$ and separately between bands $n^{\prime}=\uparrow_{x},-$ and band $m^{\prime}=\uparrow_{x},+$. Interestingly, it is possible to find a resonance with both pairs of bands fixing the light frequency such that $\hbar\omega=2|d_{z}(K)|=6\sqrt{3}t_{2}+2M$ for the right-handed wave (+). It is interesting to observe that in the present case, it is possible to fix the resonant condition just with the $K$ Dirac point.
@@
A peculiarity of this topological semimetal is that $\langle\sigma_{z}(\pi)\rangle=0$ as a result of the formation of the nodal ring semimetal around $K^{\prime}$ which is another way to justify why $\Gamma_{-}=0$ in this case. The response for one-spin polarization is now halved compared to one topological plane as a result of $\Gamma_{-}=0$. The topological semimetal in the bilayer model [22, 23] can be characterized in terms of a pair of one-half local topological markers and $\pi$-Berry phases [23] related to the two-spheres’ model associated to spin polarizations $\uparrow_{z},\downarrow_{z}$ [22]. In the one-layer graphene realization, the $\pi$ Berry phase is associated equally to the phase acquired by an electron with spin polarization $\uparrow_{x}$ in band $2$ and by an electron with spin polarization $\downarrow_{x}$ in band $1$, when circling around the $K$ Dirac. A pair of $\pi$-Berry phases is also analogous to a pair of $\pi$ winding numbers. A $\frac{1}{2}$ light response is also predicted to occur for the $\nu=\frac{1}{2}$-fractional quantum Hall effect [36].
@@
Therefore, we obtain a relation between light and topological response associated to the lowest-energy band
@@
When ${r}<3\sqrt{3}t_{2}-M$, the same light polarization would equally interact with both pairs $(n,m)$ and $(n^{\prime},m^{\prime})$ at one Dirac point measuring twice the topological invariant of the lowest band, in agreement with the quantum Hall conductivity. For ${r}>3\sqrt{3}t_{2}-M$, one pair of bands will interact with one distinct light polarization at each Dirac point. In this case, if we reproduce the steps of the calculation above we now find that $\Gamma_{+}-\Gamma_{-}=0$ at each Dirac point. In Eq. (35), the two pairs $(n,m)$ and $(n^{\prime},m^{\prime})$ are now characterized with different signs of $A_{y}(t)$ and inverted lower and upper energy eigenstates, for the same Dirac point.
We have reminded above that the light response at the $K$ and $K^{\prime}$ points for the topological Haldane model reveals the same information as the light response and therefore the quantum Hall response coming from a summation on the whole Brillouin zone [31]. This identification necessitates that the quantum Hall responses and light responses are continuous and differentiable everywhere. For the topological semimetal, this requires that the intermediate bands do not alter the result from the lowest-energy band in Sec. III. In this way, the light response from the Dirac points is identical to the quantum Hall response of the semimetal only if $\theta_{c}=\frac{\pi}{2}$. For $\theta_{c}=\frac{\pi}{2}$, we also observe that the quantum Hall response coming from $\uparrow_{x}$ particles in the second-energy band takes the form $\frac{e^{2}}{h}\widehat{C}_{\uparrow_{x},-}=-\frac{1}{2}\frac{e^{2}}{h}$ [23]. Similarly, if we now sum the topological number associated to the $\downarrow_{x}$ particles in the two energy bands, this also measures $\frac{e^{2}}{h}(\widetilde{C}_{\downarrow_{x},+}+C_{\downarrow_{x},-})=-\frac{1}{2}\frac{e^{2}}{h}$ and gives a similar interpretation to the other half-topological number. The light approach then equivalently measures the topological response $\sum_{j=\uparrow,\downarrow}C_{\alpha}$ with $|C_{\alpha}|=\frac{1}{2}$, in relation with the bilayer model [22, 18, 23].